OpenStudy (anonymous):

if * is a binary operation defined by x*y=x^2+y what is the value of 2*3

OpenStudy (anonymous):

2* 3 = 2^2 + 3 = 7

OpenStudy (anonymous):

can u please show me how u got that answer?

OpenStudy (anonymous):

substitute x and y value into the equation

OpenStudy (anonymous):

x = 2 y = 3

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

2^2 + 3 is the answer

OpenStudy (anonymous):

can u show me how you got it? im sure you are right i just have a lot of problems like these i kinda need to know how to do it